Let a and b be the roots of the equation x^2+x+1=0. The equation whose roots are a^19, b^7 is. A. x^2-x-1=0. B. x^2-x+1=0. C. x^2+x-1=0. D. x^2+x+1=0.

Grade:

2 Answers

Sher Mohammad IIT Delhi
10 years ago
a=x1, b=x2
ab=1, a+b=-1
,

=1
=-1

from above analysis , x^2+x+1 is the correct answer


Sher Mohammad
B.Tech, IIT Delhi



Nikhil Kumar
10 years ago
X^3-1=0 => (x-1)(x^2+x+1)=0 => x=1,w,w^2 totally in question we have x^2+x+1=0 means a=w, and b=w^2. ...(1) And a^19=w^19=w^18.w=w similarly b^7=w^2. Which is equal to (1) hence d